Sodium has body centred packing. If the distance between two nearest atoms is 3.7 Å, then the lattice parameter is


        (a)    3.3 Å


        (b)    3.9 Å


        (c)    4.3 Å        


            (d)    4.8 Å

Grade:10

2 Answers

Amit Saxena
10 years ago

(c)

We know that in B.C.C.

4r =  (a)

Given, 2r = 3.7 Å

2 × 3.7 × 10-10 = a

a = 4.3 × 10-10 = 4.3 Å

Amit Saxena
10 years ago

        (d)